For the reaction: 2HBr(g) + Cl2(g) -----> 2HCl(g) + Br2(I)

a) calculate delta G for this reaction using free energy values.

b) Calculate delta G at standard temperature when pressure of HBr is 4.97 atm, Cl2 is 3.56 atm and HCl is 2.18 atm.

c) Calculate Kp for this reaction.
Answer #1

For the reaction: 2HBr(g) + Cl2(g) -----> 2HCl(g) + Br2(I)

(a) calculateG for this reaction using free energy values.

4G = [24_G°(HCI) +4_Gº(Br)]-[24_Gº (HBr) - 4 Gº(C7)] ={[2x–95.27]+3. 14)+2x53.22 =-80.96 KJ/mol

(b) CalculateG at standard temperature when pressure of HBr is 4.97 atm, Cl2 is 3.56 atm and HCl is 2.18 atm.

(c) Calculate Kp for this reaction.

Kp = pHCl2 / pCl2 x pHBr2

Kp = 2.182 / 3.56 x 4.972

Kp = 4.7524 / 87.935 = 0.054

▵G = ▵Go + RTln Kp

▵Go = ▵G - RTln Kp

▵Go = -80.96 KJ - 8.314 x 298 x ln0.054

▵Go = -80.96 KJ +7.231 KJ = - 73.73 KJ